# Information-Theoretic Constraints on Variational Quantum Optimization: Efficiency Transitions and the Dynamical Lie Algebra

### v1

@@ IV Microscopic Origin
To determine the microscopic driver of this thermodynamic collapse, we analyzed the structural statistics of the operators comprising the Dynamical Lie Algebra for both topological classes. Specifically, we computed the distribution of the Pauli weights (Hamming weights) for the basis operators generated during the Lie closure process.
We observe that the macroscopic efficiency crash in the chaotic phase corresponds to a microscopic regime of maximal operator scrambling. In the Ordered ($K_{n}$) phase, the DLA operators remain sparsely supported, preserving a low average Pauli weight even as the system size scales. This sparsity implies that the information relevant to the optimization gradient remains localized in the Hilbert space, accessible to the finite-bandwidth probe of the ancilla.
@@
Figure 4: DLA Efficiency Scaling. Comparison of thermodynamic efficiency for Ordered (Complete Graph, polynomial DLA $\sim O(N^{3})$) vs. Chaotic (Spin Glass, exponential DLA $\sim O(4^{N})$) systems. The polynomial scaling of the algebra ensures sustained positive efficiency, while the exponential algebra exhibits the efficiency crash observed in Fig. 3.
@@ V Discussion
Our experimental findings suggest that the limitations of variational quantum optimization are not merely algorithmic artifacts, but manifestations of fundamental thermodynamic bounds. By treating the optimizer as a heat engine, we have identified a Carnot-like bound for quantum optimization efficiency. Analogous to how Carnot efficiency is bounded by temperature ratios, the algorithmic efficiency is bounded by the information channel capacity:
$\eta\leq\theta_{\text{gain}}\cdot\chi_{\text{Holevo}}$ (7)
where $\theta_{\text{gain}}$ is the feedback coupling strength and $\chi_{\text{Holevo}}\leq 1$ bit is the Holevo capacity of the single-ancilla channel [holevo1973bounds]. Numerically, we observe $\eta=0.456\cdot\theta_{\text{gain}}$ in the Ordered phase ($R^{2}=0.9945$), indicating operation at approximately 50% of the theoretical maximum—likely attributable to the linear response approximation inherent in gradient-based feedback and measurement backaction from the projective ancilla readout. This linear scaling with coupling strength—rather than an arbitrary empirical constant—establishes a universal limit derivable from first principles.
@@
This framework offers a physical perspective on the hardness of variational optimization. We propose that the barrier between tractable and intractable problems is observable as the transition from finite to diverging information cost. In this view, hard problems are those where the thermodynamic cost of the solution grows efficiently, requiring an exponentially increasing information flux to maintain a finite cooling rate.
While recent work by Ragone et al.[ragone2024lie] established the Dynamical Lie Algebra as the geometric predictor of trainability, the physical mechanism driving this transition remains an open question. Here, we provide evidence that the geometric ‘concentration of measure’ correlates with an information-theoretic efficiency transition, characterized by the collapse of information-to-work conversion efficiency when DLA dimension scales exponentially.
@@
Our results provide a quantum-mechanical quantification of the “Computational Irreducibility” hypothesis [wolfram2023secondlaw], demonstrating that the thermodynamic cost of optimization diverges exactly when the system dynamics become algebraically irreducible. In this framework, the Second Law of Thermodynamics arises from the limitations of a computationally bounded observer. The Carnot-like bound $\eta\leq\theta_{\text{gain}}$ (Eq. 7) can thus be interpreted as the channel capacity limit of a single-ancilla quantum observer, with the normalized efficiency $\eta/\theta_{\text{gain}}\approx 0.5$ measuring how close the optimization operates to this fundamental limit.
@@ V.1 Implications for Quantum Learning
Based on the observed scaling of the information cost, we suggest that the tractability of variational learning is constrained by the thermodynamic stability of information flow.
We define a Thermo-Efficient regime as the set of Hamiltonians for which the algorithmic efficiency remains positive ($\eta>0$) and the information susceptibility remains finite in the large-$N$ limit. Our results indicate that systems with Polynomial DLA growth (e.g., the Complete Graph $K_{n}$) belong to this regime.
Conversely, chaotic systems (e.g., Spin Glasses) with Exponential DLA growth exhibit an “Efficiency Collapse” ($\eta\to 0$). Here, the information scrambling rate exceeds the controller’s channel capacity. Our results suggest that the efficiency coefficient $\eta$ provides a practical diagnostic for variational algorithm trainability, complementary to geometric metrics like the Quantum Fisher Information [abbas2021effective]. While QFIM-based approaches require computing the full parameter Hessian, our protocol directly measures the information-work conversion capacity through a single-ancilla probe. This information bottleneck can be formalized as a violation of the Holevo-Schumacher-Westmoreland (HSW) theorem [holevo1973bounds]. The single-ancilla probe constitutes a quantum channel $\Phi$ with a classical capacity $\chi(\Phi)\leq 1$ bit. In the chaotic phase, the required information scaling $I_{\text{req}}\propto N$ exceeds this capacity.
@@ V.2 Outlook: Beyond the Barrier
Our identification of the thermodynamic limit points toward four distinct pathways to extend the tractable regime of quantum optimization:
1. Multi-Demon Bandwidth: Since the crash is driven by a channel capacity bottleneck ($I_{req}\gg 1$ bit), employing a $k$-ancilla probe could linearly increase the information bandwidth. Future work will investigate whether a “Demon Ensemble” can delay the thermodynamic crash in chaotic systems by matching the scrambling rate with parallel information extraction.
@@
3. The Temperature of Complexity: We have analyzed the zero-temperature limit of the optimizer. Introducing finite thermal noise may reveal a critical “Complexity Temperature” $T_{c}$, above which the Demon’s information gathering is erased by thermal fluctuations, establishing a hard physical bound on the operating temperature of quantum computers solving NP-hard problems.
4. The QFT Continuum Limit: Our analysis utilized discretized parameters, analogous to a Lattice Gauge Theory formulation. In the limit of infinite bit precision ($B\to\infty$), the ansatz trajectory approaches a continuous field. Future work will investigate whether the thermodynamic bounds derived here imply a fundamental computational renormalization group flow, where the tractability of finding the vacuum state of a Quantum Field Theory depends on the information scaling of its effective lattice action.
@@ VI.1 Non-Linearity Test
To establish the source of non-linear dynamics, we constructed a minimal 4-qubit circuit comparing interacting vs. non-interacting Hamiltonians. Two parameter qubits $(P_{0},P_{1})$ control rotations on two system qubits $(S_{0},S_{1})$:
@@ VI.2 Experimental Protocol: The Coherent Demon
To rigorously quantify the thermodynamic cost of optimization, we constructed a single-ancilla probe designed to isolate the information-work exchange. The protocol consists of three distinct unitary phases applied to the joint state $\rho_{SA}=\rho_{S}\otimes|0\rangle\langle 0|_{A}$:
@@
2. Correlation (Information Storage): A Hadamard gate on the ancilla converts the phase information into population differences (Z-basis). At this stage, we measure the Mutual Information $I(S:A)$ and Logarithmic Negativity $E_{N}$ to quantify the “Fuel” available for extraction. Crucially, no projective measurement is performed yet; the information is stored in quantum correlations.
3. Feedback (Actuation): We apply a Coherent Feedback operation $U_{kick}=CR_{X}(\theta_{gain})$, where the system undergoes a rotation conditioned on the ancilla state. Rather than “detecting” the gradient direction, this controlled operation creates asymmetric dynamics: amplitude at low-energy configurations receives different actuation than amplitude at high-energy configurations. The net effect is analogous to a quantum walk coin flip—one branch diffuses while the other traps.
@@ VI.4 The W-Gate Protocol
To realize the trap-diffusion mechanism, we implemented a custom “W-Gate” protocol that treats the variational parameters as quantum degrees of freedom. The key insight is that the ancilla controls which operations are applied, creating asymmetric dynamics for different energy configurations:
@@
For the Ordered phase, we utilized a Complete Graph Hamiltonian ($K_{n}$) with uniform couplings, which is known to generate a polynomially bounded DLA with dimension $O(N^{3})$ [allcock2024dynamical].
@@ VI.6 DLA Analysis
Dynamical Lie Algebra (DLA) dimensions were computed using the Lie Closure algorithm. We iteratively computed the nested commutators of the generating set $\mathcal{G}=\{iH_{S}\}\cup\{iP_{k}\}_{k=1}^{M}$, where $P_{k}$ are the Pauli generators of the ansatz, until the set closed under commutation.
@@ VI.8 Work and Information Definitions
We define the thermodynamic quantities as follows:
@@
Thermodynamic Definitions: We explicitly define the thermodynamic system boundaries to enclose only the quantum information processing degrees of freedom (qubits). Our efficiency metric $\eta$ quantifies the differential algorithmic work extracted per unit of information entropy generated within the Hilbert space, distinct from the constant macroscopic control overhead.

### v2

@@ IV Microscopic Origin
To determine the microscopic driver of this thermodynamic collapse, I analyzed the structural statistics of the operators comprising the Dynamical Lie Algebra for both topological classes. Specifically, I computed the distribution of the Pauli weights (Hamming weights) for the basis operators generated during the Lie closure process.
I observe that the macroscopic efficiency crash in the chaotic phase corresponds to a microscopic regime of maximal operator scrambling. In the Ordered ($K_{n}$) phase, the DLA operators remain sparsely supported, preserving a low average Pauli weight even as the system size scales. This sparsity implies that the information relevant to the optimization gradient remains localized in the Hilbert space, accessible to the finite-bandwidth probe of the ancilla.
@@ V Discussion
My experimental findings suggest that the limitations of variational quantum optimization are not merely algorithmic artifacts, but correlate with thermodynamic constraints. By treating the optimizer as a heat engine, I observe an empirical efficiency ceiling that scales linearly with feedback coupling strength:
$\eta\approx 0.55\cdot\theta_{\text{gain}}$ (7)
Numerically, I find this linear relationship holds with $R^{2}>0.99$ in the Ordered phase. The coefficient $\approx 0.55$ suggests the protocol operates near a robust efficiency bound, potentially related to the single-qubit ancilla’s entropy export capacity.
@@
This framework offers a physical perspective on the hardness of variational optimization. I propose that the barrier between tractable and intractable problems is observable as the transition from finite to diverging information cost. In this view, hard problems are those where the thermodynamic cost of the solution grows efficiently, requiring an exponentially increasing information flux to maintain a finite cooling rate.
While recent work by Ragone et al.[ragone2024lie] established the Dynamical Lie Algebra as the geometric predictor of trainability, the physical mechanism driving this transition remains an open question. Here, I provide evidence that the geometric ‘concentration of measure’ correlates with an information-theoretic efficiency transition, characterized by the collapse of information-to-work conversion efficiency when DLA dimension scales exponentially.
@@
My results provide a quantum-mechanical quantification of the “Computational Irreducibility” hypothesis [wolfram2023secondlaw], demonstrating that the thermodynamic cost of optimization diverges exactly when the system dynamics become algebraically irreducible. In this framework, the Second Law of Thermodynamics arises from the limitations of a computationally bounded observer. The empirical efficiency ceiling $\eta\approx 0.55\cdot\theta_{\text{gain}}$ (Eq. 7) can thus be interpreted as a channel capacity constraint of the single-ancilla probe.
@@ V.1 Implications for Quantum Learning
Based on the observed scaling of the information cost, I suggest that the tractability of variational learning is constrained by the thermodynamic stability of information flow.
I define a Thermo-Efficient regime as the set of Hamiltonians for which the algorithmic efficiency remains positive ($\eta>0$) and the information susceptibility remains finite in the large-$N$ limit. My results indicate that systems with Polynomial DLA growth (e.g., the Complete Graph $K_{n}$) belong to this regime.
Conversely, chaotic systems (e.g., Spin Glasses) with Exponential DLA growth exhibit an “Efficiency Collapse” ($\eta\to 0$). Here, the information scrambling rate exceeds the controller’s channel capacity. My results suggest that the efficiency coefficient $\eta$ provides a practical diagnostic for variational algorithm trainability, complementary to geometric metrics like the Quantum Fisher Information [abbas2021effective]. While QFIM-based approaches require computing the full parameter Hessian, my protocol directly measures the information-work conversion capacity through a single-ancilla probe. This information bottleneck may relate to the Holevo-Schumacher-Westmoreland (HSW) theorem [holevo1973bounds]. The single-ancilla probe constitutes a quantum channel $\Phi$ with a classical capacity $\chi(\Phi)\leq 1$ bit. In the chaotic phase, the required information scaling $I_{\text{req}}\propto N$ may exceed this capacity, though a formal proof of this connection is left for future work.
@@ V.2 Outlook: Beyond the Barrier
My identification of the thermodynamic limit points toward four distinct pathways to extend the tractable regime of quantum optimization:
1. Multi-Demon Bandwidth: Since the crash is driven by a channel capacity bottleneck ($I_{req}\gg 1$ bit), employing a $k$-ancilla probe linearly increases the information bandwidth. I validate this experimentally (Fig. 4): a $k=2$ ancilla ensemble extracts approximately $2\times$ the mutual information per cycle, and the total work $W=\eta\times I$ remains positive for larger system sizes compared to $k=1$. This confirms that parallel information extraction can delay the thermodynamic crash.
Figure 4: Multi-Ancilla Bandwidth Scaling. (a) Efficiency $\eta$ vs system size NN for $k=1$ (blue) and $k=2$ (green) ancillae. Non-monotonic behavior at specific NN (e.g., $N=6$) reflects known even-odd parity effects in finite-size quantum spin systems [caleca2024giant]. (b) Information bandwidth $I(S:A)$ scales approximately linearly with $k$, with measured ratio $\approx 2.0\times$. (c) Total extracted work $W=\eta\times I$ shows that $k=2$ maintains positive work extraction for larger NN than $k=1$, validating the “Multi-Demon” strategy for overcoming the channel capacity bottleneck.
@@
3. The Temperature of Complexity: I have analyzed the zero-temperature limit of the optimizer. Introducing finite thermal noise may reveal a critical “Complexity Temperature” $T_{c}$, above which the Demon’s information gathering is erased by thermal fluctuations, establishing a hard physical bound on the operating temperature of quantum computers solving NP-hard problems.
4. The QFT Continuum Limit: My analysis utilized discretized parameters, analogous to a Lattice Gauge Theory formulation. In the limit of infinite bit precision ($B\to\infty$), the ansatz trajectory approaches a continuous field. Future work will investigate whether the thermodynamic bounds derived here imply a fundamental computational renormalization group flow, where the tractability of finding the vacuum state of a Quantum Field Theory depends on the information scaling of its effective lattice action.
@@ VI.1 Non-Linearity Test
To establish the source of non-linear dynamics, I constructed a minimal 4-qubit circuit comparing interacting vs. non-interacting Hamiltonians. Two parameter qubits $(P_{0},P_{1})$ control rotations on two system qubits $(S_{0},S_{1})$:
@@ VI.2 Experimental Protocol: The Coherent Demon
To rigorously quantify the thermodynamic cost of optimization, I constructed a single-ancilla probe designed to isolate the information-work exchange. The protocol consists of three distinct unitary phases applied to the joint state $\rho_{SA}=\rho_{S}\otimes|0\rangle\langle 0|_{A}$:
@@
2. Correlation (Information Storage): A Hadamard gate on the ancilla converts the phase information into population differences (Z-basis). At this stage, I measure the Mutual Information $I(S:A)$ and Logarithmic Negativity $E_{N}$ to quantify the “Fuel” available for extraction. Crucially, no projective measurement is performed yet; the information is stored in quantum correlations.
3. Feedback (Actuation): I apply a Coherent Feedback operation $U_{kick}=CR_{X}(\theta_{gain})$, where the system undergoes a rotation conditioned on the ancilla state. Rather than “detecting” the gradient direction, this controlled operation creates asymmetric dynamics: amplitude at low-energy configurations receives different actuation than amplitude at high-energy configurations. The net effect is analogous to a quantum walk coin flip—one branch diffuses while the other traps.
@@ VI.4 The W-Gate Protocol
To realize the trap-diffusion mechanism, I implemented a custom “W-Gate” protocol that treats the variational parameters as quantum degrees of freedom. The key insight is that the ancilla controls which operations are applied, creating asymmetric dynamics for different energy configurations:
@@
For the Ordered phase, I utilized a Complete Graph Hamiltonian ($K_{n}$) with uniform couplings, which is known to generate a polynomially bounded DLA with dimension $O(N^{3})$ [allcock2024dynamical].
@@ VI.6 DLA Analysis
Dynamical Lie Algebra (DLA) dimensions were computed using the Lie Closure algorithm. I iteratively computed the nested commutators of the generating set $\mathcal{G}=\{iH_{S}\}\cup\{iP_{k}\}_{k=1}^{M}$, where $P_{k}$ are the Pauli generators of the ansatz, until the set closed under commutation.
@@ VI.8 Work and Information Definitions
I define the thermodynamic quantities as follows:
@@
Thermodynamic Definitions: I explicitly define the thermodynamic system boundaries to enclose only the quantum information processing degrees of freedom (qubits). My efficiency metric $\eta$ quantifies the differential algorithmic work extracted per unit of information entropy generated within the Hilbert space, distinct from the constant macroscopic control overhead.